# TRiSM for Agentic AI: A Review of Trust, Risk, and Security Management in LLM-based Agentic Multi-Agent Systems

### v1

@@ Abstract
Agentic AI systems, built on large language models (LLMs) and deployed in multi-agent configurations, are redefining intelligent autonomy, collaboration and decision-making across enterprise and societal domains. This review presents a structured analysis of Trust, Risk, and Security Management (TRiSM) in the context of LLM-based agentic multi-agent systems (AMAS). We begin by examining the conceptual foundations of agentic AI, its architectural differences from traditional AI agents, and the emerging system designs that enable scalable, tool-using autonomy. The TRiSM in the agentic AI framework is then detailed through four pillars governance, explainability, ModelOps, and privacy/security each contextualized for agentic LLMs. We identify unique threat vectors and introduce a comprehensive risk taxonomy for the agentic AI applications, supported by case studies illustrating real-world vulnerabilities. Furthermore, the paper also surveys trust-building mechanisms, transparency and oversight techniques, and state-of-the-art explainability strategies in distributed LLM agent systems. Additionally, metrics for evaluating trust, interpretability, and human-centered performance are reviewed alongside open benchmarking challenges. Security and privacy are addressed through encryption, adversarial defense, and compliance with evolving AI regulations. The paper concludes with a roadmap for responsible agentic AI, proposing research directions to align emerging multi-agent systems with robust TRiSM principles for safe, accountable, and transparent deployment.
00footnotetext: *Shaina Raza and Ranjan Sapkota contributed equally and are corresponding authors.
Emails: shaina.raza@torontomu.ca, rs2672@cornell.edu, mk2684@cornell.edu, c.emmanouilidis@rug.nl.
Keywords Agentic AI, LLM-based Multi-Agent Systems, TRISM, AI Governance, Explainability, ModelOps, Application Security, Model Privacy, Autonomous Agents, Trustworthy AI, Risk Management, AI Safety, Privacy-Preserving AI, Adversarial Robustness, Human-in-the-Loop
@@ 1.1 Background and Motivation
The global market for AI agents is projected to grow from \$5.4 billion in 2024 to \$7.6 billion in 2025  miller2025aiagents . By mid-2025, over 70% of enterprise AI deployments are expected to involve multi-agent or action-based systems, reflecting a dramatic shift from traditional single-agent or rule-based conversational models lyzr2025state . An AI agent, typically defined as a computational entity that perceives its environment and takes actions to achieve goals  russell1997rationality , has undergone a fundamental transformation. Whereas early agents were task-specific and deterministic, modern systems, referred to as Agentic AI, now integrate large language models (LLMs), tools, and persistent memory to support complex planning and coordination sapkota2025ai ; de2023emergent .
@@
From isolated agents to Agentic AI. The autonomous software agents date back to the 1990s castelfranchi1998modelling , the advent of LLM-driven multi-agent ecosystems yang2024multi since 2023 has made “Agentic AI” a distinctly new and rapidly expanding research focus sapkota2025ai . Traditional agents automated narrow tasks, such as information retrieval chiu2007biomimetic ; lew2006content , data summarization tanasa2005advanced ; devault2014simsensei ; hao2012mri , and dialogue response griol2013automatic . Driven by single step logic genesereth2012logical or scripted rules, they lacked deep reasoning, adaptability, and persistence spronck2006adaptive ; szafir2012pay .
@@
In contrast, Agentic AI consists of collaborative agents with specialized roles (e.g., planner, coder, analyst), enabled by LLMs and tool use acharya2025agentic . These systems dynamically decompose tasks, share context, and pursue high-level goals across long timelines qian2023chatdev ; wu2023autogen ; hong2023metagpt . This shift represents not just a technological upgrade but a qualitative leap in complexity and autonomy, marking the emergence of machine collectives capable of emergent, decentralized behavior.
Trust, risk, and security challenges. However, this evolution introduces serious challenges. Unlike deterministic agents zeigler2014object , Agentic AI systems can produce non-linear, opaque decisions, increasing the risks of failure, bias, and unintended consequences. For example, a multi-agent supply chain optimizer may autonomously coordinate between procurement and logistics agents, yet inadvertently leak sensitive information or violate compliance protocols if safeguards are absent lim2012memory .
@@
Traditional evaluation and safety frameworks, built for static or single-function AI , are no longer sufficient. This underscores the urgent need for a new paradigm that integrates trust, risk, and security as core design principles. To address this gap, recent frameworks like AI TRiSM (Trust, Risk, and Security Management) IBM2025AITRiSM ; habbal2024artificial propose lifecycle-level controls, including explainability, secure model orchestration, and privacy management. These are essential for deploying agentic systems in high-stakes domains such as finance, healthcare, and defense.
@@
Problem Statement Agentic AI systems now coordinate multiple LLM-powered agents, yet no unified framework exists for managing their trust, risk, and security (TRiSM). High-profile lapses, e.g., an autonomous research agent citing unverifiable medical sources or a forecasting swarm drifting off-target due to memory inconsistencies show that traditional safeguards do not scale to dynamically evolving, multi-agent workflows. A structured TRiSM perspective is therefore crucial to mitigate orchestration failures, collusion, data leakage, and opaque decision paths. This review distills current findings and outlines practical guidelines to build secure, reliable, and trustworthy Agentic AI ecosystems.
@@
Scope and Objectives This survey maps the TRiSM (Trust, Risk, and Security Management) landscape for LLM-powered, multi-agent “agentic” AI. We frame four pillars explainability, security, lifecycle governance, and privacy and show why safeguards designed for bounded, rule-based agents cannot contain open-world, self-orchestrating collectives sapkota2025ai ; acharya2025agentic . Our goal is to give researchers and practitioners a concise, unified reference for building and governing trustworthy Agentic AI systems.
@@
Necessity of this Survey Agentic AI research often concentrates on agent modelling or task coordination and rarely tackles adversarial threats, lifecycle governance, or large-scale explainability. As deployments reach high-stakes arenas healthcare, science, finance the absence of a unified TRiSM framework exposes stakeholders to opaque reasoning and unmanaged risks. This survey closes that gap by synthesising TRiSM principles for LLM-driven multi-agent systems and providing actionable guidance for researchers, engineers, and policy makers. A comparison with related surveys will be presented in Table 1.
@@ 1.2 Contributions
In this survey, we address the urgent need for a unified governance framework for LLM-based Agentic AI systems. Our framework is shown in Figure 1. Our key contributions are as follows:
  • A Conceptual TRISM Framework Tailored to Agentic AI We introduce a structured Trust, Risk, and Security Management (TRISM) framework specifically designed to accommodate the unique characteristics of multi-agent, LLM-driven systems. This framework defines four pillars: Explainability, ModelOps, Application Security, and Model Privacy, and explains how each pillar addresses the challenges of autonomy, coordination, and persistent memory in Agentic AI.
  • A Threat Taxonomy for LLM-Powered Agents We develop a detailed taxonomy of risks and threat vectors that arise when multiple LLM agents collaborate. By categorizing vulnerabilities, such as prompt injection, memory poisoning, collusive failures, and emergent misbehavior, we clarify how attack surfaces expand in agentic environments and why traditional LLM security measures are insufficient.
  • A Comparative Evaluation of Existing Methods We survey and compare state-of-the-art techniques for each TRISM pillar as applied to Agentic AI. For example, we examine how explainable-AI methods (e.g., LIME, SHAP, decision provenance) have been adapted for multi-agent workflows, how ModelOps best practices must be extended to manage multiple LLM instances, and which security/hardening measures (e.g., prompt hygiene, sandboxing, access controls) are most effective.
  • Future Research Directions for Trustworthy, Secure Agentic AI Building on our TRISM analysis, we outline concrete directions for next-generation research and setup roadmap items aim to guide researchers and practitioners toward robust, scalable, and compliant deployments of Agentic AI.
This study provide a unified reference for researchers, engineers, and policymakers who seek to design, evaluate, and govern LLM-based Agentic AI systems in high-stakes domains.
@@ 1.3 Paper Organization
The remainder of this paper is structured as follows. In Section 2, we present our literature-search methodology, including databases queried, inclusion criteria, and classification strategies. Section 3 reviews the fundamental characteristics and typical architectures of Agentic AI systems, establishing the groundwork for subsequent TRISM analysis. Section 4 introduces the TRISM framework and its four core pillars : Explainability, ModelOps, Application Security, and Model Privacy. Building on this foundation, Section 5 identifies the primary risks, vulnerabilities, and attack surfaces inherent in multi-agent AI ecosystems. In Section 6, we explore trust-building strategies and specialized explainability techniques tailored to Agentic AI, while Section 7 details advanced security mechanisms and privacy-preserving methods. Section 8 synthesizes our findings and outlines future research directions and policy recommendations. Section 9 concludes the paper. A list of key terms used throughout this paper is provided in Table 9.
1.4 Related Work
Table 1: Comparison of Related Surveys on LLM-based Multi-Agent Systems and TRiSM Aspects
Survey Adversarial Threats Lifecycle Governance Large-Scale Explainability TRiSM Integration LLM-Specific Application Domains Actionable Guidance
Guo et al. (2024) guo2024largelanguagemodelbased (simulated env’ts) ~ (research challenges)
Chen et al. (2025) chen2025surveyllmbasedmultiagentsystem ~ (task/simulation focus) ~ (future research areas)
Yan et al. (2025) yan2025selftalkcommunicationcentricsurveyllmbased (mentions diverse) ~ (future work directions)
Tran et al. (2025) tran2025multiagentcollaborationmechanismssurvey (networks, QA, etc.) ~ (open challenges)
Lin et al. (2025) lin2025creativityllmbasedmultiagentsystems ~ (creative tasks) ~ (roadmap for research)
Fang et al. (2025) fang2025trustworthyaisafetybias ~ ~ (health, finance cited) ~ (technical “next steps”)
This Survey (2025) (high-stakes domains) (practical guidance)
@@ 1.4 Related Work
LLM-Based Multi-Agent Surveys (Technical Focus): Prior surveys on LLM-driven multi-agent systems primarily emphasize system architectures, agent capabilities, and domain-specific applications of agentic AI, but often overlook critical TRiSM considerations. For instance, existing reviews guo2024largelanguagemodelbased ; chen2025surveyllmbasedmultiagentsystem focus on simulated environments, inter-agent communication, and performance benchmarks, yet fail to address adversarial threats, lifecycle governance, or system-wide explainability. Other targeted surveys, such as yan2025selftalkcommunicationcentricsurveyllmbased on natural language communication and tran2025multiagentcollaborationmechanismssurvey on collaboration mechanisms, explore interaction protocols and use cases in domains like QA systems, 5G, and Industry 5.0. While the former briefly acknowledges security concerns and the latter highlights deployment contexts, both narrowly center on coordination strategies without examining robustness, governance structures, or explainability at scale.
Trustworthy/Responsible AI Surveys (Broad Trust Focus): Surveys in the TRiSM domain address ethical and risk-related concerns such as alignment, fairness, and privacy attacks fang2025trustworthyaisafetybias , but typically overlook explainability, governance, and multi-agent dynamics. While they offer actionable recommendations for general ML systems, they do not account for LLM-based coordination in high-stakes domains.
Current Survey (LLM Multi-Agent + TRiSM in High-Stakes Domains): This survey uniquely integrates TRiSM principles into the analysis of LLM-based multi-agent systems, specifically targeting high-stakes domains such as healthcare, science, and finance. In contrast to earlier surveys (shown in Table 1) that underemphasize governance and security, this work foregrounds adversarial threats (e.g., vulnerabilities in agent interaction), end-to-end lifecycle governance (from training data to deployment oversight), and large-scale explainability. These concerns are treated as integral to system design, not peripheral. Additionally, this survey provides actionable recommendations for researchers, engineers, and policymakers, bridging gaps left by both technical and responsible-AI surveys.

### v2

@@ Abstract
Agentic AI systems, built upon large language models (LLMs) and deployed in multi-agent configurations, are redefining intelligence, autonomy, collaboration, and decision-making across enterprise and societal domains. This review presents a structured analysis of Trust, Risk, and Security Management (TRiSM) in the context of LLM-based Agentic Multi-Agent Systems (AMAS). We begin by examining the conceptual foundations of Agentic AI and highlight its architectural distinctions from traditional AI agents. We then adapt and extend the AI TRiSM framework for Agentic AI, structured around four key pillars: Governance, Explainability, ModelOps, and Privacy/Security , each contextualized to the challenges of multi-agent LLM systems. A novel risk taxonomy is proposed to capture the unique threats and vulnerabilities of Agentic AI, ranging from coordination failures to prompt-based adversarial manipulation. To support practical assessment in Agentic AI works, we introduce two novel metrics: the Component Synergy Score (CSS), which quantifies the quality of inter-agent collaboration, and the Tool Utilization Efficacy (TUE), which evaluates the efficiency of tool use within agent workflows. We further discuss strategies for improving explainability in Agentic AI , as well as approaches to enhancing security and privacy through encryption, adversarial robustness, and regulatory compliance. The review concludes with a research roadmap for the responsible development and deployment of Agentic AI, outlining critical directions to align emerging systems with TRiSM principles for safe, transparent, and accountable operation.
keywords:
Agentic AI, LLM-based Multi-Agent Systems, TRiSM, AI Governance, Explainability, ModelOps, Application Security, Model Privacy, AI Agents, Trustworthy AI, Risk Management, AI Safety, Privacy-Preserving AI, Adversarial Robustness, Human-in-the-Loop
@@ 1.1 Background and Motivation
The global market for Artificial Intelligence (AI) agents is projected to grow from \$5.4 billion in 2024 to \$7.6 billion in 2025  [1]. By mid-2025, over 70% of enterprise AI deployments are expected to involve multi-agent or action-based systems, reflecting a dramatic shift from traditional single-agent or rule-based conversational models [2]. An AI agent, typically defined as a computational entity that perceives its environment and takes actions to achieve goals  [3], has undergone a fundamental transformation. While early agents were task-specific and deterministic, modern systems have evolved significantly ,increasing operational autonomy by over 70% on benchmark tasks [4]. These Agentic AI systems now integrate large language models (LLMs), external tools, and persistent memory to support complex planning and coordinated decision-making [5, 6].
This rapid evolution presents new opportunities but also surfaces significant challenges in coordination, control, transparency, and risk assessment across heterogeneous agents. In particular, the shift toward Agentic AI architectures amplifies and evolves longstanding concerns around safety, fairness, interpretability, and systemic harms. While such issues have been central to machine learning (ML) systems broadly, they become particularly critical in Agentic AI, where autonomous agents operate in high-stakes domains like healthcare, law, and public services, areas where their decisions can carry profound and lasting societal consequences.
@@
From isolated agents to Agentic AI The autonomous software agents date back to the 1990s [7], however, the advent of LLM-driven multi-agent ecosystems [8] since 2023 has made “ Agentic AI ” a distinctly new and rapidly expanding research focus [5]. Traditional agents primarily automated narrow tasks, such as information retrieval [9, 10], data summarization [11], and dialogue response [12]. Typically driven by single step logic [13] or scripted rules, these agents lacked the deep reasoning, adaptability, and persistence required for complex, multi-step problem solving [14, 15].
@@
In contrast, Agentic AI consists of collaborative agents with specialized roles (e.g., planner, coder, analyst), enabled by LLMs and tool use [16]. These systems dynamically decompose tasks, share context, and pursue high-level goals across long timelines [17, 18, 19]. This shift reflects not only a change in technology but also a fundamental increase in complexity and autonomy, marking the emergence of machine collectives capable of exhibiting emergent and decentralized behavior.
1.2 Trust, Risk, and Security Challenges in Agentic AI
The rise of Agentic AI introduces complex challenges that extend beyond those of traditional ML systems. Unlike deterministic agents [20], LLM-based multi-agent systems exhibit emergent and often opaque decision-making, increasing the risk of cascading errors, biases, and unintended behaviors. These characteristics necessitate a robust approach to have trust and ensuring trustworthiness, in high-stakes domains such as healthcare, finance, and law.
Trust and Trustworthiness
In the context of Agentic AI, trust refers to a user willingness to rely on an AI system, while trustworthiness denotes whether the system consistently behaves in a safe, fair, and predictable manner, thereby deserving that trust. This distinction is critical in high-stakes settings, where undue reliance on an untrustworthy system can lead to severe consequences, such as misdiagnoses in healthcare or biased rulings in legal applications. The stochastic nature of LLM reasoning [21] further complicates trustworthiness, as randomness in output generation, while promoting creativity and ensemble diversity, introduces inconsistency and non-repeatability. These factors hinder traceability, verification, and compliance auditing, all of which are essential for trustworthy systems.
Security Risks in Multi-Agent Systems
Security risks escalate as autonomous agents gain access to tools, external APIs, and persistent memory, amplifying the potential for privacy breaches, adversarial misuse, or regulatory violations in the absence of robust human oversight [22]. For example, a multi-agent supply chain optimizer may autonomously coordinate between procurement and logistics agents but inadvertently leak sensitive data or violate compliance protocols (e.g., GDPR or HIPAA) if safeguards are inadequate. The distributed and cooperative nature of multi-agent systems introduces unique vulnerabilities that traditional security measures often fail to address.
Specific Threats: Prompt Injection, Spoofing, and Impersonation
LLM-based agents are susceptible to novel security threats exploiting their language-based interfaces and cooperative behaviors. One prominent threat is prompt injection [23], where attackers craft inputs with hidden or malicious instructions to manipulate agent behavior. Recent studies [24] describe “prompt infection,” a phenomenon where malicious prompts propagate from one agent to another, akin to a computer virus spreading across networks [25]. Such corrupted outputs can lead to cascading effects, including data leaks, fraudulent transactions, misinformation, or coordinated misbehavior within an agent society.
Another critical vulnerability involves spoofing and impersonation [26]. Spoofing entails faking an entity’s identity or credentials (e.g., using a false name or address) to deceive a system, while impersonation involves actively pretending to be a specific agent or user to gain trust, access, or privileges. In multi-agent systems, where agents rely on credentials or tokens for authentication during coordination, an adversary could steal an API key or mimic a trusted peer to issue unauthorized commands or access sensitive information.
Privacy Risks in Multi-Agent LLM Systems
LLM-based agents often need to handle user data, proprietary business information, or other sensitive inputs to fulfill their tasks [27]. In a multi-agent context, this challenge is amplified by the fact that agents may share information with each other (e.g. via a shared memory store or message passing) to collaborate.
Limitations of Traditional Frameworks
Traditional evaluation and safety frameworks, designed for deterministic agents or conventional AI systems, are increasingly inadequate for addressing the autonomy and complexity of Agentic AI. These frameworks often fail to account for emergent behaviors, inter-agent dependencies, or the dynamic nature of LLM-driven systems, leaving gaps in security, trust, and risk management.
Towards a TRiSM Framework for Agentic AI
@@ 1.2 Trust, Risk, and Security Challenges in Agentic AI
To address these gaps, recent frameworks like AI TRiSM (Trust, Risk, and Security Management) [28, 29] propose lifecycle-level controls, including explainability, secure model orchestration, and privacy management. These components are vital for deploying Agentic AI in high-stakes domains such as finance, healthcare, and defense, where trust, security, and accountability are non-negotiable.
Figure 1: Taxonomy of TriSM for Agentic AI systems presented in this review. It includes five major sections: Fundamentals of Agentic AI Systems ( Agentic AI vs. Traditional AI Agents, LLM-Based Architecture), The TRISM Framework (Governance, Explainability, ModelOps, Application Security, Privacy), Threats and Risks , Trust and Explainability, and Evaluation Metrics (including Human-Centered and Composite Metrics).
1.3 Problem Statement
@@
Agentic AI systems increasingly rely on the coordination of multiple LLM-powered agents to perform complex, long-horizon tasks. However, no unified framework exists for managing their Trust, Risk, and Security Management (AI TRiSM) across dynamic, multi-agent workflows. Recent failures, such as, an autonomous research agent citing unverifiable medical sources or collaborative swarms generating inconsistent forecasts due to memory drift, highlight that traditional safeguards are inadequate [30]. These traditional systems are vulnerable to orchestration breakdowns, agent collusion, data leakage, and opaque decision-making pathways. To ensure safety and reliability in Agentic AI systems, a structured TRiSM perspective is urgently needed. This study summarizes emerging findings and offers actionable guidelines to build secure, transparent, and trustworthy Agentic AI ecosystems.
1.4 Scope and Objectives
@@
Objectives This study is a review article that maps the TRiSM (Trust, Risk, and Security Management) [28] landscape as it applies to LLM-powered, multi-agent “Agentic AI” systems. We structure the discussion around four foundational pillars of TRiSM [31]: explainability, security, lifecycle governance, and privacy. We argue that conventional safeguards developed for bounded, rule-based agents are inadequate for managing the complexities of open-ended, self-directed agent collectives. The objective of this review is to offer researchers and practitioners a coherent and actionable reference for the design, deployment, and governance of trustworthy Agentic AI ecosystems.
For the sake of this study, we define Agentic AI as Multi-Agent Systems (MAS) [32] powered by LLMs that exhibit autonomous planning, tool use, and memory retention, with or without human oversight. In contrast to classical MAS, these systems exhibit emergent reasoning capabilities and dynamically adapt their workflows based on interactions with tools, users, and other agents.
Scope This paper is intended as a conceptual survey and synthesis. It does not propose a new Agentic AI architecture or introduce synthetic benchmark evaluations. Rather, it integrates emerging perspectives on trust, risk, and security to support future research and responsible development in this evolving area.
@@
Necessity of this Survey Agentic AI research has largely centered on agent modeling, planning, and task coordination, while comparatively overlooking critical dimensions such as adversarial robustness, lifecycle governance, and system-level explainability. As these systems are increasingly deployed in high-stakes domains - such as healthcare, scientific research, and finance—the absence of an integrated TRiSM framework exposes stakeholders to opaque decision-making and unmanaged risks. This survey addresses that gap by synthesizing TRiSM principles in the context of LLM-driven multi-agent systems and providing actionable guidance for researchers, engineers, and policymakers.
To the best of our knowledge, few existing surveys systematically explore LLM-based multi-agent systems through the lens of trust, risk, and security. Table 1 underscores this gap by comparing prior reviews with respect to their coverage of TRiSM dimensions.
@@ 1.5 Contributions
Our key contributions in this review are:
  1. 1.
    We propose the first TRiSM framework that is explicitly tailored to systems in which several LLM-powered agents collaborate and cooperate. The framework is organized around four pillars: Explainability, Lifecycle Governance, Application Security, and Model Privacy. It targets the specific governance challenges created by autonomous decision making, inter-agent coordination, and persistent memory (see Fig. 1 and Sec. 5).
  2. 2.
    We offer a systematic taxonomy of risks that emerge when multiple agents interact (Sec. 6). This taxonomy discusses prompt injection, memory poisoning, collusive failure, emergent misbehavior, and related threats, and it shows why defenses designed for single-model LLM applications are not sufficient.
  3. 3.
    We introduce a five-dimension evaluation framework that covers Trustworthiness, Explainability, User-centered Performance, Coordination, and Composite Assessment (Sec. 7.3). To support this template, we define two new metrics: the Component Synergy Score (CSS), which measures the quality of inter-agent collaboration, and the Tool Utilization Efficacy (TUE), which measures how correctly and efficiently agents invoke external tools.
  4. 4.
    We review state-of-the-art techniques for each TRiSM pillar (Secs. 78) and examine methods such as Local Interpretable Model-Agnostic Explanations (LIME), SHapley Additive exPlanations (SHAP), decision provenance, prompt hygiene, sandboxing, and ModelOps (model operations) pipelines. These techniques must be critically re-evaluated and adapted to suit environments involving multiple interacting LLM agents.
  5. 5.
    Building on the identified research gaps, we outline a research agenda that prioritizes scalable, verifiable, and regulation-compliant Agentic AI (Sec. 9).
@@ 1.6 Paper Organization
The remainder of this paper is structured as follows. Section 2 discuss the related work. Section 3 presents the literature search methodology. Section 4 reviews the core characteristics and architectures of Agentic AI systems, providing a foundation for the TRiSM analysis. Section 5 introduces the TRiSM framework and its four pillars: Explainability, ModelOps, Application Security, and Model Privacy. Section 6 outlines key risks, vulnerabilities, and attack surfaces in multi-agent AI ecosystems. Section 7 discusses trust-building and explainability techniques, while Section 8 covers security and privacy-enhancing methods. Section 9 synthesizes insights and proposes directions for research and policy. Section 10 concludes the paper with final reflections on the implications of TRiSM for the responsible development of Agentic AI. A list of key terms used throughout this paper is provided in Table A.1.
2 Related Work
Table 1: Comparison of Related Surveys on LLM-Based Multi-Agent Systems and TRiSM Aspects. Existing work offers limited coverage of trust, risk, and security in Agentic AI.
[…]
Guo et al. (2024) [33] ✓ (simulated env’ts) ~ (research challenges)
Chen et al. (2025) [34] ~ (task/simulation focus) ~ (future research areas)
Yan et al. (2025) [35] ✓ (mentions diverse) ~ (future work directions)
Tran et al. (2025) [36] ✓ (networks, QA, etc.) ~ (open challenges)
Lin et al. (2025) [37] ~ (creative tasks) ~ (roadmap for research)
Fang et al. (2025) [38] ~ ~ ✓ (health, finance cited) ~ (technical “next steps”)
This Survey (2025) ✓ (high-stakes domains) ✓ (tailored guidance)
@@ 2 Related Work
[…]
We present the related work in this section.
2.1 LLM-Based Multi-Agent Surveys (Technical Focus)
Prior surveys on LLM-driven multi-agent systems predominantly focus on technical dimensions such as system architectures, agent capabilities, and domain-specific applications of Agentic AI, yet they rarely address critical TRiSM (Trust, Risk, and Security Management) considerations. For example, reviews like [33] and [34] delve into simulated environments, inter-agent communication protocols, and performance benchmarks. These works excel at detailing how agents coordinate or scale but fall short in examining adversarial threats, such as prompt injection attacks or agent collusion, lifecycle governance (e.g., continuous monitoring or regulatory compliance), or system-wide explainability, especially for emergent behaviors in complex systems.
More specialized surveys, such as [35] on natural language communication and [36] on collaboration mechanisms, explore interaction protocols and practical use cases in domains like question-answering systems, 5G networks, and Industry 5.0. While [35] briefly notes security risks tied to communication channels, and [36] mentions deployment contexts, both remain narrowly focused on coordination strategies. They lack deeper analysis of robustness against adversarial manipulation (e.g., Byzantine failures), governance frameworks for multi-agent deployments, or explainability challenges at scale.
2.2 Trustworthy/Responsible AI Surveys (Broad Trust Focus)
Surveys addressing TRiSM in AI, such as [38], tackle ethical and risk-related issues including alignment with human values, fairness in decision-making, and defenses against privacy attacks (e.g., data leakage). However, these works typically target general ML systems [39] and overlook the unique dynamics of LLM-based multi-agent systems. For instance, while they may advocate for privacy-preserving techniques like differential privacy, they do not address privacy risks in inter-agent data sharing or the distributed nature of LLM-driven systems. Moreover, critical TRiSM aspects [40] like governance (e.g., auditing autonomous agents) and explainability (e.g., tracing decisions across multiple agents) are often absent. As a result, their recommendations, while valuable for single-agent or centralized AI, do not fully translate to the complexities of multi-agent coordination in high-stakes domains.
2.3 Current Study (LLM Multi-Agent)
This review distinguishes itself by weaving TRiSM principles: adversarial robustness, end-to-end lifecycle governance, and system-wide explainability, into the study of LLM-based multi-agent systems, with a specific lens on high-stakes domains like healthcare, science, and finance. Unlike prior studies (see Table 1), which offer limited insight into governance and security, this work positions these elements as foundational design priorities. For example, it addresses targeted threats like backdoor attacks in distributed LLMs and governance challenges such as ensuring compliance with regulations (e.g., HIPAA in healthcare or GDPR in finance) across autonomous agents - areas largely ignored by technical surveys. This work provides actionable guidance for researchers, engineers, and policymakers, and fills critical gaps in both the technical multi-agent literature and broader responsible AI landscape.